Lean meat is marinated and added moisture through brining. In brining, a mixture of salt and water is used to soak the meat for a few hours or days before cooking.

Basic Brine: Make 3880 grams brine for 2.72 kg of meat

Procedure:

a. Mix 250 grams of salt (NaCl) with 3880 grams of water.

b. If you make sweet brine, mix 250 grams of sugar (C12H22O11) to solution of salt and water.

Problem: What is the molal concentration of the brine solution (250 g salt + 3880 g water)?


1
Expert's answer
2022-01-31T17:15:05-0500

First find the mass in 1000g solute.

"=\frac{1000×250}{3880}=64.43g"

Molality is the number of moles per 1000g solvent.

"=\frac{64.43}{23+35.5}=1.1"m